i have two coordinates in the x axis being(1 and 4) and in the y axis (3 and 2) how do i find the distance between two cordinates?
Answer by venugopalramana(3286)

i have two coordinates in the x axis being(1 and 4) and in the y axis (3 and 2) how do i find the distance between two cordinates?
do you mean the points are (1,3)and (4,2)...if so
d= sqrt{(x2-x1)^2+(y2-y1)^2}=sqrt{(4-1)^2+(2-3)^2}=sqrt(9+1)=sqrt(10)